OpenStudy (anonymous):

A wooden pyramid, 6 inches tall, has a square base. A carpenter increases the dimensions of the wooden pyramid by a factor of 3 and makes a larger pyramid with the new dimensions. Describe in complete sentences the ratio of the volumes of the two pyramids.

OpenStudy (campbell_st):

the ratio of the linear measurements will be 1:3 the height of 6inches will become 18 in the new pyramid. which volumes... the linear measurements will be cubed to get the ratio of the volumes \[1^3 : 3^3\] or 1: 27 so the larger pyramid with measurements 3 time the smaller will be 27 times larger. e.g. if the pyramid is 6 inches tall and the square base has side length 1 inch V = 1/3 x 1^2 x 6 V = 2 cubic inches larger.... height = 18 base side length 3 V = 1/3 x 3^2 x 18 V = 54 cubic inches ratio of the volumes is 2:54 or 1:27

OpenStudy (anonymous):

thanks